OpenStudy (anonymous):

simplify -9[12-8(-3)*(-5)+1]

OpenStudy (anonymous):

-9(12+24*-5 + 1) = -9(12 - 120 + 1) = -9 * -107 = 963

OpenStudy (anonymous):

963